OpenStudy (anonymous):

solve the intequality below for x 3/4x + 2 _<3x-1

OpenStudy (anonymous):

3/4 _< x

OpenStudy (anonymous):

so sorry 4/3 _< x

OpenStudy (zarkon):

is this the problem \[\frac{3}{4}x + 2\le3x-1\]

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

\[x\ge \frac{4}{3}\]

OpenStudy (anonymous):

3/4 x + 2 - 3x <= -1 - 2 1/4 x <= -3 x >= -3 * - 4/9 x >= 4/3

OpenStudy (anonymous):

thanks